How long will it take for the population of a certain country to double if its annual growth rate is 3.5%? (Round to the nearest year.) Use the exponential growth model P(t) = P e kt.
0
1 year
9 years
57 years
20 years

Answer by Alan3354(21550)

How long will it take for the population of a certain country to double if its annual growth rate is 3.5%? (Round to the nearest year.) Use the exponential growth model P(t) = P e kt.
--------------
Future P = P*e^kt
2*P = P*e^0.035t
2 = e^0.035t
ln(2) = 0.035t
t = ln(2)/0.035
t = 19.8 years ~ 20 years